OpenStudy (anonymous):

Find an equation with real coefficients of a polynomial function f that has the given characteristics. Degree: 3 Zeros: 1, 2-3i y-intercept: -26

jimthompson5910 (jim_thompson5910):

x = 1, x = 2-3i, x = 2+3i x - 1 = 0, x - (2 -3i) = 0, x - (2 + 3i) = 0 a(x-1)( x - (2 -3i) )( x - (2 + 3i) ) = 0 a(x-1)( x - 2 +3i )( x - 2 - 3i ) = 0 a(x-1)( (x - 2) +3i )( (x - 2) - 3i ) = 0 a(x-1)( (x - 2)^2 -(3i)^2 ) = 0 a(x-1)( x^2-4x+4+9 ) = 0 a(x-1)( x^2-4x+13) = 0 a[x( x^2-4x+13 )-1( x^2-4x+13 )] = 0 a[ x^3-4x^2+13x-x^2+4x-13] = 0 a[ x^3-5x^2+17x-13] = 0 ax^3-5ax^2+17ax-13a = 0 So the function is f(x) = ax^3-5ax^2+17ax-13a But we know that the y-intercept is (0, -26) So f(0) = -26 This means that a(0)^3-5a(0)^2+17a(0)-13a = -26 -13a = -26 a = -26/(-13) a = 2 So the function is now f(x) = ax^3-5ax^2+17ax-13a f(x) = 2x^3-5(2)x^2+17(2)x-13(2) f(x) = 2x^3-10x^2+34x-26 So the answer is f(x) = 2x^3-10x^2+34x-26